Factor completely. Show all work necessary.
4x^2-9

Answer by richwmiller(17219)

This we know as a difference in squares
4x^2 is one square and 9 is the other
3^2=9
(2x)^2=4x^2
The middle terms cancel each other out with plus and minus.
a^2-b^2=(a-b)*(a+b)
(2x-3)*(2x+3)